OpenStudy (anonymous):

Can someone please help me to solve this?w^2/3=2w^1/3+8 Thank you!

OpenStudy (anonymous):

nobody wanna help me?! (:

OpenStudy (anonymous):

Put the apples on one side and the oranges on the other.

OpenStudy (anonymous):

2w^1/3-w^2/3=-8?

OpenStudy (anonymous):

Subtract 2-1 and subtract (1/3)-(2/3)

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

\[w ^{-\frac{ 1 }{ 3 }}=-8\]

OpenStudy (anonymous):

now?

OpenStudy (anonymous):

Maybe we should have worked it another way to avoid a negative in the exponent.\[w ^{\frac{ 2 }{ 3 }}-2w ^{\frac{ 1 }{ 3 }}=8\]

OpenStudy (anonymous):

you don't even know what to do, you are just pretending like you know